Income Statement
Sales/Revenue
Total revenue 76,480,000,000
Cost of Revenue 25,110,000,000
Gross Profit 51,370,000,000
Operating Exp 31,810,000,000
Selling General & Admin -
Other Operating Expense 3,390,000,000
Unusual Expense 16,180,000,000
EBIT after Unusual Expense 2,130,000,000
Non Operating Income/Expense 385,000,000
Equity in Affiliates (Pretax) 1,020,000,000
Interest Expense 17,670,000,000
Pretax Income 16,370,000,000
Other After Tax Income (Expense) 1,300,000,000
Consolidated Net Income -
Minority Interest Expense 1,300,000,000
Net Income -
Discontinued Operations 1,300,000,000
Net Income After Extraordinaries -
Preferred Dividends 1,300,000,000
Net Income Available to Common

480,000,000

Balance Sheet
Assets
Cash & Short Term Investments $ 18,300,000,000.00
Total Accounts Receivable $ 13,490,000,000.00
Inventories $     8,770,000,000.00
Other Current Assets $     2,540,000,000.00
Total Current Assets $ 43,090,000,000.00
Net Property, Plant & Equipment $ 17,010,000,000.00
Total Investments and Advances $        751,000,000.00
Long-Term Note Receivable -
Intangible Assets $ 85,130,000,000.00
Other Assets $     4,220,000,000.00
Total Assets $157,300,000,000.00
Liabilities & Shareholders' Equity
Accounts Payable $     7,310,000,000.00
Income Tax Payable $     1,850,000,000.00
Other Current Liabilities $ 17,470,000,000.00
Total Current Liabilities $ 30,540,000,000.00
Long-Term Debt $ 30,680,000,000.00
Stockholders' Equity
Other Liabilities $ 17,490,000,000.00
Total Liabilities $ 97,140,000,000.00
Common Equity (Total) $ 60,160,000,000.00
Total Shareholders' Equity $ 60,160,000,000.00
Total Equity $ 60,160,000,000.00
Liabilities & Shareholders' Equity $157,300,000,000.00
Current Ratio $               1.41
Inv Turnover $               8.72
Total asset turnover $               0.49
Debt Ratio $               0.19
Debt to Equity ratio $               1.61

6.  Basic earning power

7. ROE

8. P/E Ratio

9. Market to Book Ratio

10.  The Dupont Equation:

We decompose the ROE into three ratios:

ROE = (Net profit margin) (Total asset utilization)(Equity Multiplier)
Answer #1

6)Basic earning power= EBIT/total assets
=2130000000/157300000000=1.35%
7)ROE=net income/equity
=480000000/60160000000=0.80%
8)P/E ratio= price per share/earning per share
=
9)Market to book ratio= price per share/book value per share
=

8 and 9 can not be found since market price per share not given
10)ROE=(net profit margin)*(total asset util)*(equity multiplier)
=(net profit/sales)*(sales/assets)*(assets/equity)
=(480000000/76480000000)*(76480000000/157300000000)*(157300000000/60160000000)
=(0.63%)*(0.49)*(2.61)
=0.80%